OpenStudy (anonymous):

a^2b^3c = 256/27 find the min value of a+b+c

OpenStudy (anonymous):

Take the numbers as a/2,a/2,b/3,b/3,b/3,c apply AM>= GM you will get the answer

OpenStudy (anonymous):

Let us say a,b,c are three numbers in AP then AM = (a+b+c) /3 and if a,b,c are three numbers in GP GM = (a*b*c)^1/3

OpenStudy (anonymous):

does it make sense to you?